OpenStudy (anonymous):

A curve y=3 sqrt x and the line y=x interesct at O and P. Find the coordinates of P

OpenStudy (anonymous):

It's simultaneous equation right? But, how do I solve it? x= 3 sqrt x

OpenStudy (anonymous):

right, sqrt x=3 so, x=9 y=x=9 so, co-ordinates of P=(9,9)

OpenStudy (lgbasallote):

x^2 = 9x x^2-9x = 0 x(x-9) = 0 x = 0 x = 9 did i do it right? haha

OpenStudy (anonymous):

9,9 assuming O is origin.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Wait, there are different answers.... How do I do this?

OpenStudy (lgbasallote):

or 0,0 assuming that O is not origin...it's not really specified which of O and P is origin

OpenStudy (anonymous):

x^2=9 right? so isn't that x=3 or x=-3?

OpenStudy (anonymous):

O is the origin

OpenStudy (anonymous):

x^2=9x

OpenStudy (lgbasallote):

x = 3 sqrt x square that.. x^2 = 3^2 (sqrt x)^2 so you still have an x

OpenStudy (anonymous):

So, what's the answer? -3 and 3, or what?

OpenStudy (lgbasallote):

either (0,0) or (9,-9)

OpenStudy (lgbasallote):

(9.9) i mean

OpenStudy (anonymous):

Are you sure?

OpenStudy (lgbasallote):

considering 3 of us answered that..i am assuming we are sure

OpenStudy (callisto):

\[y=3\sqrt{x}\]Put y=x into the equation \[x=3\sqrt{x}\]Square both sides\[x^2=(3\sqrt{x})^2\]\[x^2=9x\]\[x^2-9x=0\]\[x(x-9)=0\]x=0 or x-9 =0 x=0 or x=9 therefore (0,0) and (9,9) are the intercepts (since y=x)

OpenStudy (callisto):

As O is the origin, that is (0,0), P is (9,9)

OpenStudy (anonymous):

Thanks, that cleared it up :)

OpenStudy (anonymous):